OpenStudy (anonymous):

Select the symbol below that fills in the blank to correctly complete the mathematical statement: 4/7____ 8/14 > < =

OpenStudy (anonymous):

In fractions, you can multiply both the numerator and the denominator by the same number and get the same number. So, for example, 1/2 = 1*2 / 2*2 = 2/4. Then: 4/7 = 4*2 / 7*2 = 8/14. So the symbol is "=".

jimthompson5910 (jim_thompson5910):

4/7 ? 8/14 4*14 ? 7*8 56 ? 56 56 = 56 So 4/7 = 8/14

OpenStudy (anonymous):

Its equal because we can divide the second term 8/14 by 2 and we will get 4/7 which is equal to the left side.